OpenStudy (anonymous):

Find the area of a parallelogram with sides of 12 inches and 8 inches if one of the angles is 120

OpenStudy (anonymous):

I remember this question

OpenStudy (anonymous):



OpenStudy (anonymous):

work out the height using trigonometry and its 12 * height

OpenStudy (amistre64):

\[{{12\ cos(120)}\choose{0\ sin(120)}}^{0}_{12sin(120)}=12sin(120)-0=12sin(120)\]

OpenStudy (amistre64):

drop the negative for area tho ;)

OpenStudy (anonymous):

tanx its for geometry 4√3.is the anwer

OpenStudy (amistre64):

0 1 2 3 4 0 30 45 60 90 sin(120) = -sqrt(3)/2 if my memory device serves me

OpenStudy (amistre64):

you sure it aint 6sqrt(3) for the area?

OpenStudy (amistre64):

i forgot an 8 lol ... hmm

OpenStudy (amistre64):

height = 8sin(60) in the drawing base times height = area 12 * 8sin(60) = 96sqrt(3)/2